# Proper scoring rules for estimation and forecast evaluation

### v2

@@ 3.2 Local scoring rules
Arguably, the most popular and important of all proper scoring rules is the logarithmic score given by $S_{\mathrm{log}}(\mathbb{P},\mathbf{y})=-\log p(\mathbf{y})$, compare Examples 5 and 9. It is a local scoring rule of order 00, and remarkably, the only differentiable scoring rule of this kind (Bernardo, 1979). Empirical risk minimization for the logarithmic score is equivalent to maximum likelihood estimation which is the foundation of classical parametric inference. The unparalleled historical success of maximum likelihood is in great part because it furnishes closed-form estimates for many important parametric families of distributions and is asymptotically optimal under rather mild regularity conditions.
@@
Theorem 12 can be used to derive a simple proof that the logarithmic score is the only differentiable local scoring rule of order 00. Indeed, let $S(\mathbb{P},\mathbf{y})=s(\mathbf{y},p(\mathbf{y}))$ be such a scoring function. The function $s(\mathbf{y},z)$ is differentiable in $z$, and its entropy is $H(\mathbb{P})=\int s(\mathbf{x},p(\mathbf{x}))p(\mathbf{x})\,\mathrm{d}\mathbf{x}$. By Equation (7) and using Equation (9), we must have
@@
where $p$ is the density of $\mathbb{P}$, see Examples 6 and 10. It is a local scoring rule of order $2$. The value of the Hyvärinen score does not change if we replace $p$ by $c\cdot p$ for some constant $c>0$, which implies that the score can be evaluated even with non-normalized $p$.
@@
A fairly general characterization of local scoring rules of arbitrary order has been obtained by Parry et al. (2012) using a calculus of variations approach. In particular, there are no local scoring rules of odd orders, and all local scoring rules of even order $\geq 2$ can be computed without knowledge of the normalization constant of the density like in the case of the Hyvärinen score. A more explicit characterization for the particular case of order $2$ was derived by Ehm and Gneiting (2012) using an approach based on Theorem 12. For example, Ehm and Gneiting (2012, Example 5.3) introduce the log-cosh score
@@ 3.3 $f$-Scores, $gf$-scores and generalized kernel scores
Another twist of the idea is to replace the integral in Equation 17 with a different entropy, for example, the entropy of a kernel score as at Equation 11, that is, considering entropies of the form
@@ 3.4 Proper scoring rules induced by elicitable functionals
for all $z\in\mathcal{Z}$, $\mathbb{P}\in\mathcal{P}$, where $Y\sim\mathbb{P}$, subject to the existence of expectations, with equality if and only if $z=T(\mathbb{P})$. Such a scoring function $s$ is called strictly consistent for the functional $T$. Relevant functionals that are elicitable are the mean, quantiles, expectiles, and ratios of expections. Some important functionals like the variance are not elicitable. There is a large literature on elicitability with possible starting points being Lambert et al. (2008); Gneiting (2011); Fissler and Ziegel (2016).
@@ 4 Scoring Rule Decompositions
The familiar bias-variance decomposition of the squared error loss can be generalized to Bregman divergences (Pfau, 2013). Since proper scoring rules are extensions of Bregman loss functions, we can derive a similar decomposition of the performance of a probabilistic predictor. Following the conditional setting in Section 1.1, let $\smash{\mathbb{P}_{\hat{\theta},x}}$ be an estimator of $\mathbb{P}_{Y|X=x}$ where the parameter $\smash{\hat{\theta}}=\smash{\hat{\theta}(\mathcal{D})}$ has been estimated from a sample $\mathcal{D}$ of $(X,Y)$. Then we can write
@@
where ¯$\bar{\mathbb{P}}_{x}=\operatorname*{arg\,min}_{\mathbb{P}}\mathbb{E}[d(\smash{\mathbb{P}_{\hat{\theta},x}},\mathbb{P})]$ and $\mathbb{P}_{Y|X=x}=\operatorname*{arg\,min}_{\mathbb{P}}\mathbb{E}[S(\mathbb{P},Y)\,|\,X=x]$ is the best predictor, assuming the two exist. If we think of ¯$\smash{\bar{\mathbb{P}}_{x}}$ as the “expectation” of the predictor $\smash{\mathbb{P}_{\hat{\theta},x}}$ over $\mathcal{D}$, then just like the original bias-variance decomposition, the variance term measures how much the estimator $\smash{\mathbb{P}_{\hat{\theta},x}}$ varies around ¯$\smash{\bar{\mathbb{P}}_{x}}$ and the bias term measures the distance between ¯$\smash{\bar{\mathbb{P}}_{x}}$ and the best predictor $\mathbb{P}_{Y|X=x}$. Indeed, the two decompositions coincide for the Brier score. While ¯$\bar{\mathbb{P}}_{x}$ corresponds to a left Bregman projection, commonly studied in optimization, the optimal predictor $\mathbb{P}_{Y|X=x}$ is a right Bregman projection, which is of greater importance in statistics (Bauschke et al., 2018).
@@
where $\mathbb{P}_{\smash{\raisebox{-3.61664pt}{$\hat{}\,$}}}$ denotes the conditional law of the outcome $Y$ given the forecast $\hat{\mathbb{P}}$ and $\bar{\mathbb{P}}$ is the marginal law of $Y$, often called the climatological forecast in meteorology. The miscalibration (MCB) term measures how different the forecast $\hat{\mathbb{P}}$ is from the optimal calibrated forecast $\mathbb{P}_{\smash{\raisebox{-3.61664pt}{$\hat{},$}}}$ with the same information, which is the conditional law of $Y$ given $\hat{\mathbb{P}}$. Discrimination ability (DSC) quantifies how much more descriminative the optimal calibrated forecast $\mathbb{P}_{\smash{\raisebox{-3.61664pt}{$\hat{}\,$}}}$ for the given information is in comparison to only marginal information about $Y$. Finally, the uncertainty term (UNC) informs about the general difficulty of the forecasting problem.
@@ 5.1.2 Transformation Models
$\hat{S}(\mathbb{P}_{\theta,\mathbf{x}},y)=\frac{1}{2m(m-1)}\sum_{i=1}^{m}\sum_{i^{\prime}:i^{\prime}\neq i}h(g_{\theta}(\mathbf{x},\mathbf{N}_{i}),g_{\theta}(\mathbf{x},\mathbf{N}_{i^{\prime}}))-\frac{1}{m}\sum_{i=1}^{m}h(g_{\theta}(\mathbf{x},\mathbf{N}_{i}),y)$
@@ 5.2.2 Weighted Scoring Rules
In many statistical applications such as risk management (Holzmann and Klar, 2017) and high impact weather prediction (Allen et al., 2023), certain outcomes are of considerably greater consequence to the user than others and naturally, we want the scoring rule to weigh forecast performance for such outcomes accordingly. To achieve this, Allen et al. (2023) propose using kernel scores with kernels modified using thresholding-weighting or vertical-rescaling, that is, $h^{\prime}(x,y)=h(v(x),v(y))$ and $h^{\prime}(x,y)=h(x,y)w(x)w(y)$, respectively, by using a chaining function $v:\mathcal{Y}\to\mathcal{Y}$ or weighing function $w:\mathcal{Y}\to\mathbb{R}$. By choosing $h$ to be a translation invariant kernel, we can control the weight exclusively through the chaining and weighting functions. This generalizes the threshold-weighted CRPS introduced by Gneiting and Ranjan (2011), which also introduced a quantile-weighted version of the CRPS. Towards the same end, Diks et al. (2011) propose weighted versions of the logarithmic score based on which Holzmann and Klar (2017) propose a different approach to modifying scoring rules, which generally does not lead to kernel scores even if the original score was a kernel score.
@@ Acknowledgements
The authors would like to thank an anonymous reviewer, Sam Allen, Philip Dawid, Ritabrata Dutta, Sebastian Engelke, Rafael Frongillo, Tilmann Gneiting, Pierre Jacob, Alexander Jordan, Kristof Kraus, Andrea Nava, Romain Pic, and Xinwei Shen for comments, inspiring and fruitful discussions, and useful pointers to the literature.
@@ Appendix A Proofs for Section 2
$\lim_{\alpha\to 0^{+}}\frac{1}{\alpha^{2}}\Big{[}H((1-\alpha)\mathbb{P}+\alpha\mathbb{Q})-H(\mathbb{P})-\alpha\langle\nabla_{\mathbb{P}}H,\mathbb{Q}\rangle\Big{]}\\ =\frac{1}{2}\langle\nabla^{2}_{\mathbb{P}}H(\mathbb{Q}),\mathbb{Q}\rangle=\frac{1}{2}\iint\nabla^{2}_{\mathbb{P}}H(x,x^{\prime})\,\mathrm{d}\mathbb{Q}(x)\,\mathrm{d}\mathbb{Q}(x^{\prime}).$ (18)
@@
where $\{\nu_{r}\}$ are probability measures on $\mathbb{S}^{d-1}$ for $r\in(0,\infty)$ and $\rho$ is a $\sigma$-finite measure on $(0,\infty)$. In order for $\int_{\mathbb{R}^{d}}\min(1,\|\mathbf{x}\|^{2})\,\mathrm{d}\mu(\mathbf{x})<\infty$ to hold, we must have $\int_{0}^{\infty}\min(1,r^{2})\,\mathrm{d}\rho(r)<\infty$.
@@
    which implies $\rho((0,cr))=\frac{1}{c^{\alpha}}\rho((0,r))$ for $c,r>0$. By fixing $r=1$ and replacing $c$ with $r$, we get $\rho((0,r))=\frac{1}{r^{\alpha}}\rho((0,1))$, which implies that $\rho$ is absolutely continuous with respect to the Lebesgue measure on $(0,\infty)$ and

### v3

@@ 3.2 Local scoring rules
Arguably, the most popular and important of all proper scoring rules is the logarithmic score given by $S_{\mathrm{log}}(\mathbb{P},\mathbf{y})=-\log p(\mathbf{y})$, compare Examples 5 and 9. It is a local scoring rule of order 0, and remarkably, the only differentiable scoring rule of this kind (Bernardo, 1979). Empirical risk minimization for the logarithmic score is equivalent to maximum likelihood estimation which is the foundation of classical parametric inference. The unparalleled historical success of maximum likelihood is in great part because it furnishes closed-form estimates for many important parametric families of distributions and is asymptotically optimal under rather mild regularity conditions.
@@
Theorem 12 can be used to derive a simple proof that the logarithmic score is the only differentiable local scoring rule of order 0. Indeed, let $S(\mathbb{P},\mathbf{y})=s(\mathbf{y},p(\mathbf{y}))$ be such a scoring function. The function $s(\mathbf{y},z)$ is differentiable in $z$, and its entropy is $H(\mathbb{P})=\int s(\mathbf{x},p(\mathbf{x}))p(\mathbf{x})\,\mathrm{d}\mathbf{x}$. By Equation (7) and using Equation (9), we must have
@@
where $p$ is the density of $\mathbb{P}$, see Examples 6 and 10. It is a local scoring rule of order 222. The value of the Hyvärinen score does not change if we replace $p$ by $c\cdot p$ for some constant $c>0$, which implies that the score can be evaluated even with non-normalized $p$.
@@
A fairly general characterization of local scoring rules of arbitrary order has been obtained by Parry et al. (2012) using a calculus of variations approach. In particular, there are no local scoring rules of odd orders, and all local scoring rules of even order 2$\geq$ 2≥ 2 can be computed without knowledge of the normalization constant of the density like in the case of the Hyvärinen score. A more explicit characterization for the particular case of order 222 was derived by Ehm and Gneiting (2012) using an approach based on Theorem 12. For example, Ehm and Gneiting (2012, Example 5.3) introduce the log-cosh score
@@ 3.3 $f$-Scores, $gf$-scores and generalized kernel scores
Another twist on the idea is to replace the integral in Equation 17 with a different entropy, for example, the entropy of a kernel score as at Equation 11, that is, considering entropies of the form
@@ 3.4 Proper scoring rules induced by elicitable functionals
for all $z\in\mathcal{Z}$, $\mathbb{P}\in\mathcal{P}$, where $Y\sim\mathbb{P}$, subject to the existence of expectations, with equality if and only if $z=T(\mathbb{P})$. Such a scoring function $s$ is called strictly consistent for the functional $T$. Relevant functionals that are elicitable are the mean, quantiles, expectiles, and ratios of expectations. Some important functionals like the variance are not elicitable. There is a large literature on elicitability with possible starting points being Lambert et al. (2008); Gneiting (2011); Fissler and Ziegel (2016).
@@ 4 Scoring Rule Decompositions
[…]
@@
where ¯$\bar{\mathbb{P}}_{x}=\operatorname*{arg\,min}_{\mathbb{P}}\mathbb{E}[d(\smash{\mathbb{P}_{\hat{\theta},x}},\mathbb{P})]$ and $\mathbb{P}_{Y|X=x}=\operatorname*{arg\,min}_{\mathbb{P}}\mathbb{E}[S(\mathbb{P},Y)\,|\,X=x]$ is the best predictor, assuming the two exist. If we think of ¯$\smash{\bar{\mathbb{P}}_{x}}$ as the “expectation” of the predictor $\smash{\mathbb{P}_{\hat{\theta},x}}$ over $\mathcal{D}$, then just like in the original bias-variance decomposition, the variance term measures how much the estimator $\smash{\mathbb{P}_{\hat{\theta},x}}$ varies around ¯$\smash{\bar{\mathbb{P}}_{x}}$ and the bias term measures the distance between ¯$\smash{\bar{\mathbb{P}}_{x}}$ and the best predictor $\mathbb{P}_{Y|X=x}$. Indeed, the two decompositions coincide for the Brier score. While ¯$\bar{\mathbb{P}}_{x}$ corresponds to a left Bregman projection, commonly studied in optimization, the optimal predictor $\mathbb{P}_{Y|X=x}$ is a right Bregman projection, which is of greater importance in statistics (Bauschke et al., 2018).
@@
where $\mathbb{P}_{\smash{\raisebox{-3.61664pt}{$\hat{}\,$}}}$ denotes the conditional law of the outcome $Y$ given the forecast $\hat{\mathbb{P}}$ and $\bar{\mathbb{P}}$ is the marginal law of $Y$, often called the climatological forecast in meteorology. The miscalibration (MCB) term measures how different the forecast $\hat{\mathbb{P}}$ is from the optimal calibrated forecast $\mathbb{P}_{\smash{\raisebox{-3.61664pt}{$\hat{},$}}}$ with the same information, which is the conditional law of $Y$ given $\hat{\mathbb{P}}$. Discrimination ability (DSC) quantifies how much more discriminative the optimal calibrated forecast $\mathbb{P}_{\smash{\raisebox{-3.61664pt}{$\hat{}\,$}}}$ for the given information is in comparison to only marginal information about $Y$. Finally, the uncertainty term (UNC) informs about the general difficulty of the forecasting problem.
@@ 5.1.2 Transformation Models
$\hat{S}(\mathbb{P}_{\theta,\mathbf{x}},y)=\frac{1}{m}\sum_{i=1}^{m}h(g_{\theta}(\mathbf{x},\mathbf{N}_{i}),y)-\frac{1}{2m(m-1)}\sum_{i=1}^{m}\sum_{i^{\prime}:i^{\prime}\neq i}h(g_{\theta}(\mathbf{x},\mathbf{N}_{i}),g_{\theta}(\mathbf{x},\mathbf{N}_{i^{\prime}}))$
@@ 5.2.2 Weighted Scoring Rules
In many statistical applications such as risk management (Holzmann and Klar, 2017) and high impact weather prediction (Allen et al., 2023), certain outcomes are of considerably greater consequence to the user than others and naturally, we want the scoring rule to weigh forecast performance for such outcomes accordingly. To achieve this, Allen et al. (2023) propose using kernel scores with kernels modified using threshold-weighting or vertical-rescaling, that is, $h^{\prime}(x,y)=h(v(x),v(y))$ and $h^{\prime}(x,y)=h(x,y)w(x)w(y)$, respectively, by using a chaining function $v:\mathcal{Y}\to\mathcal{Y}$ or weighting function $w:\mathcal{Y}\to\mathbb{R}$. By choosing $h$ to be a translation invariant kernel, we can control the weight exclusively through the chaining and weighting functions. This generalizes the threshold-weighted CRPS introduced by Gneiting and Ranjan (2011), which also introduced a quantile-weighted version of the CRPS. Towards the same end, Diks et al. (2011) propose weighted versions of the logarithmic score based on which Holzmann and Klar (2017) propose a different approach to modifying scoring rules, which generally does not lead to kernel scores even if the original score was a kernel score.
@@ Acknowledgements
The authors would like to thank an anonymous reviewer, Sam Allen, Philip Dawid, Ritabrata Dutta, Sebastian Engelke, Rafael Frongillo, Tilmann Gneiting, Daniel Hug, Pierre Jacob, Alexander Jordan, Kristof Kraus, Andrea Nava, Romain Pic, Mathias Trabs, and Xinwei Shen for comments, inspiring and fruitful discussions, and useful pointers to the literature.
@@ Appendix A Proofs for Section 2
$\lim_{\alpha\to 0^{+}}\frac{1}{\alpha^{2}}\Big{[}H((1-\alpha)\mathbb{P}+\alpha\mathbb{Q})-H(\mathbb{P})-\alpha\langle\nabla_{\mathbb{P}}H,\mathbb{Q}-\mathbb{P}\rangle\Big{]}\\ =\frac{1}{2}\langle\nabla^{2}_{\mathbb{P}}H(\mathbb{Q}-\mathbb{P}),\mathbb{Q}-\mathbb{P}\rangle\\ =\frac{1}{2}\iint\nabla^{2}_{\mathbb{P}}H(x,x^{\prime})\,\mathrm{d}(\mathbb{Q}-\mathbb{P})(x)\,\mathrm{d}(\mathbb{Q}-\mathbb{P})(x^{\prime}).$ (18)
@@
where $\{\nu_{r}\}$ are probability measures on $\mathbb{S}^{d-1}$ for $r\in(0,\infty)$ and $\rho$ is a $\sigma$-finite measure on (0,)(0$\infty$)( 0 , ∞ ). In order for $\int_{\mathbb{R}^{d}}\min(1,\|\mathbf{x}\|^{2})\,\mathrm{d}\mu(\mathbf{x})<\infty$ to hold, we must have $\int_{0}^{\infty}\min(1,r^{2})\,\mathrm{d}\rho(r)<\infty$.
@@
    which implies $\rho((0,cr))=\frac{1}{c^{\alpha}}\rho((0,r))$ for $c,r>0$. By fixing $r=1$ and replacing $c$ with $r$, we get $\rho((0,r))=\frac{1}{r^{\alpha}}\rho((0,1))$, which implies that $\rho$ is absolutely continuous with respect to the Lebesgue measure on (0,)(0$\infty$)( 0 , ∞ ) and